Mimi is one third as old as George. In 5 years she will be one half as old as George. How old is Mimi?
Answer by stanbon(75887)

Mimi is one third as old as George. In 5 years she will be one half as old as George. How old is Mimi?
-----
Equations:
m = (1/3)g
(m+5) = (1/2)(g+5)
----
Rearrange:
g = 3m
2(m+5) = g+5
----
Substitute for "g" and solve for m"
2m+10 = 3m+5
m = 5 (Mimi age now)
====================
Cheers,
Stan H.